# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v5

@@ 3.4.1 The proof
First, we assume that two of the face axes at $\mathbf{{S}}_{0}$ are distinct. By Lemma 3.3, the alternative is that all six face axes are the same. This case will be handled using other methods in [ACG+].
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming that the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sect. 3.4.5 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
Figure 6 depicts the completed cube configuration that we must construct. There are 6 axes corresponding to condition (ii), one for each face. Each face axis is the common line of the 4 ring planes on the edges of the face. As mentioned above, we restrict our attention to the first case in Lemma 3.3, whereby the 3 axes of the complete faces of the 7-configuration are not all identical and have a common point $O$, as shown in Figure 6. The dual plane $o$ is taken as the plane at infinity in this figure, thus the cones of ring contact are all cylinders through the conics $\mathbf{{R}}^{k}_{j}$ where $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{k}$ make ring contact. The 4 ring points on a face lie on six distinct spears in $o$.
@@
Under the assumption that two of the faces axes at $\mathbf{{S}}_{0}$ are distinct, it suffices to construct a unique quadric $\mathbf{{T}}^{0}$ such that: for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
@@
The first equality follows from the fact that the intersection points of $\mathbf{{S}}_{i}$ with $\Lambda^{0n}_{ij}$ necessarily occur on $\mathbf{{R}}_{i}$ since as point sets $\Lambda^{0n}_{ij}\subset p^{0}_{i}$, the second equality follows from the definition (10) of $\Lambda^{0n}_{ij}$, the third from the ring contact of $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{n}$ in the plane $p^{n}_{i}$ and the rest of the equalities follow by symmetrical reasoning. Both $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ are in contact with $\mathbf{{T}}^{n}$ at these intersection points, so slicing the common tangent planes with $p$ yields the lines of double contact of $\mathbf{{R}}_{i}$ and $\mathbf{{R}}_{j}$. Therefore, under the assumption that $P$ and $p$ are not incident, we obtain the conclusion that the three conics $\mathbf{{R}}_{i}$ are mutually in double contact with chords of contact $\Lambda^{0n}_{ij}$, however these are concurrent in the point $O$, so by Lemma 2.5 must be in the same double contact pencil which would imply that the 3 axes $\Lambda^{0n}_{ij}$ coincide which we’ve already seen cannot be the case. Therefore $P$ is incident to $p$. ∎
@@
To complete the cube diagram and prove the theorem, we apply Lemma 3.4 to obtain a unique quadric containing all three conics that is in ring contact with $\mathbf{{S}}_{i}$ along $\mathbf{{R}}^{0}_{i}$ with ring point $P^{0}_{i}$ for $i=1,2,3$.
Using this fact, we now provide a synthetic proof of the 8-quadric theorem, followed by a self-contained coordinate-based proof which includes a proof of Lemma (3.4) as part of the construction.
Recall that to complete the proof we must show that for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $F^{0j}_{km}$ have a common axis.
@@
We now show that condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.3, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.3. ∎
@@ 3.4.3 Refinement of Penrose 8-quadric theorem
The above discussion contains a proof of the following refinement of Thm. 3.2.
Theorem 3.3.
Given a regular Penrose 7-quadric configuration. Let $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ be the quadric that, by Thm. 3.2, completes the configuration. Consider the 3 face axes and 3 face spears of the cube faces meeting at the eighth vertex. Then:
@@ 3.4.4 Proof of Extrusion Lemma
We now turn to the promised proof of Lemma 3.1: Given a regular Penrose 7-conic configuration, it is possible to extrude it to a regular Penrose 7-quadric configuration.
@@ 3.4.5 Proof of the 8-quadric theorem with a common contact element
3.5 Examples of 8-quadric configurations
The aim of this section is to illustrate the structure of the 8-quadric theorem by constructing several examples step by step and, by slicing them, to obtain all the famous 2D-porisms mentioned in the Introduction. Although the 8-quadric theorem was stated for regular quadrics, much of the reasoning in Sect. 3.4 can be traced through degenerate cases such as those in this section. We provide some indications of necessary modifications in the logic of the proof in order to apply to these examples; however, we leave the delineation of the general conditions for allowable degeneracies, and the complete enumeration of the possible cube configurations that may result, to a future work dedicated solely to that subject.
@@ 3.5 Examples of 8-quadric configurations
Consult Fig. 7. We start with a sphere (green) as quadric $\mathbf{{S}}_{0}$. Then we choose 3 planes $p_{0}^{i}$ intersecting the sphere in 3 (blue) rings $\mathbf{{R}}_{0}^{i}$ such that each pair of rings meet in a pair of points (white). They form a conical octahedron, featured above in Lemma 3.4. The corresponding cones touching the sphere along those rings are taken as quadrics $\mathbf{{T}}^{i}$ (yellow). Now for each pair of cones the intersection $\mathbf{{T}}^{i}\cap\mathbf{{T}}^{j}$ consists of two conics (red). These six conics meet three at a time in the 8 yellow points, the common intersection $\mathbf{{T}}^{1}\cap\mathbf{{T}}^{2}\cap\mathbf{{T}}^{3}$ of all the three cones. We call this configuration a conical cube, whereby the yellow points represent the vertices and each red conic represents a pair of opposite edges of the cube.
@@
In the latter case the final quadric $\mathbf{{T}}^{0}$ is regular (directly given by Lemma 3.4). Since each quadric $\mathbf{{S}}_{i}$ degenerates to a conic, all 3 rings $\mathbf{{R}}_{i}^{k}$ on it also collapse to this conic and each axis $\Lambda_{ij}^{0k}$ is the intersection of two ring planes. In the former case, on the other hand, the final quadric $\mathbf{{T}}^{0}$ is a point pair. This case, however, cannot as easily be subsumed into the logic of section 3.4 since the final 3 axes are identical: $\Lambda^{01}_{23}=\Lambda^{02}_{13}=\Lambda^{03}_{12}$. To picture this in terms of a limit, we consider a sequence of shrinking ellipsoids, the $\mathbf{{S}}_{i}$ getting flatter and $\mathbf{{T}}^{0}$ getting thinner and finally collapsing to a line segment connecting the two points, all planes through either point ultimately being tangent planes enveloping the ellipsoid in the limit.
@@
Figure 8: Left: Three concurrent conics of the conical cube induce a hexagon circumscribing a circle in the plane $p$ through the ring points of the cones and thus proves the Brianchon theorem. Right: In the case of three non-concurrent conics the 6 points split into two triangles. This figure proves the Poncelet theorem.
@@
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guaranteed by Lemma 3.4. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@
Consider 3 planes $p_{0}^{i}$ tangent to a quadric $\mathbf{{S}}_{0}$ at points $P_{0}^{i}$. In order to see the rings $\mathbf{{R}}_{0}^{i}$ as real objects we choose here a regulus as quadric $\mathbf{{S}}_{0}$, thus each ring is a line pair contained in the regulus. Consult Fig. 10. This is the case referred to as X-contact in section 2.2, whereby the ring point is incident to the ring plane. We want to picture these line pairs as ‘quadrics’ $\mathbf{{T}}^{i}$, although the situation is more complex. They are in fact double planes when considered point-wise and double points when considered plane-wise, but in an X-contact pencil we can identify them with the line pairs themselves. We can imagine these ‘quadrics’ as the limit of a sequence of reguli all containing the line pair $\mathbf{{S}}_{0}\cap p_{0}^{i}$, as they get flatter and flatter.
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 3.1 and 3.3) or by pairwise envelopes (Example 3.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
@@ 4 Algebraic proof of the Penrose theorems
We now turn to a purely algebraic proof of Penrose’s 8-conic theorem, which presents explicit parametrized formulas for all 8 conics of a Penrose cube, starting from $\mathbf{{S}}_{0}$ and its 3 chords of contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$, and $\mathbf{{T}}^{3}$, respectively. It presents these formulas without detailed derivations. It focuses on the case of point-wise conics; similar formulas exist also for line-wise conics.
4.1 The proof
The proof makes repeated use of the fact that two point-wise conics are in double contact with each other if and only if their pencil contains a double line. That is, there exists $a$ and $b$ such that
@@ 4.1 The proof
where ${m}$ is the joining line of the contact points. When we have established this relation for a pair of conics, we have proved that the two conics are in double contact.
We start with a conic $\mathbf{{S}}_{0}$ and a line $\mathbf{p}$. Consider the sum:
@@
Then $\mathbf{{T}}^{1}$ and $\mathbf{{S}}_{0}$ are in double contact because we have $\mathbf{{T}}^{1}-\mathbf{{S}}_{0}=\mathbf{p}^{2}$, the square of a line as stated above.
@@
There is now the important property that the conic $\mathbf{{S}}_{0}$ can move along two conics in double contact with it, while staying in double contact with both. The family of such conics can be expressed using $\mathbf{{S}}_{0}$, the chords of contact $\mathbf{q}$ and $\mathbf{r}$, and a free parameter $a$:

### v7

@@ 4.4.1 The proof
First, we assume that two of the face axes at $\mathbf{{S}}_{0}$ are distinct. By Lemma 12, the alternative is that all six face axes are the same. This case is handled in Appendix A.
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming that the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sect. 4.4.5 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 10.
Figure 6 depicts the completed cube configuration that we must construct. There are 6 axes corresponding to condition (ii), one for each face. Each face axis is the common line of the 4 ring planes on the edges of the face. As mentioned above, we restrict our attention to the first case in Lemma 12, whereby the 3 axes of the complete faces of the 7-configuration are not all identical and have a common point $O$, as shown in Figure 6. The dual plane $o$ is taken as the plane at infinity in this figure, thus the cones of ring contact are all cylinders through the conics $\mathbf{{R}}^{k}_{j}$ where $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{k}$ make ring contact. The 4 ring points on a face lie on six distinct spears in $o$.
@@
Under the assumption that two of the faces axes at $\mathbf{{S}}_{0}$ are distinct, it suffices to construct a unique quadric $\mathbf{{T}}^{0}$ such that: for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $\mathcal{F}^{0j}_{km}$ have a common axis.
@@
The first equality follows from the fact that the intersection points of $\mathbf{{S}}_{i}$ with $\Lambda^{0n}_{ij}$ necessarily occur on $\mathbf{{R}}_{i}$ since as point sets $\Lambda^{0n}_{ij}\subset p^{0}_{i}$, the second equality follows from the definition (10) of $\Lambda^{0n}_{ij}$, the third from the ring contact of $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{n}$ in the plane $p^{n}_{i}$ and the rest of the equalities follow by symmetrical reasoning. Both $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ are in contact with $\mathbf{{T}}^{n}$ at these intersection points, so slicing the common tangent planes with $p$ yields the lines of double contact of $\mathbf{{R}}_{i}$ and $\mathbf{{R}}_{j}$. Therefore, under the assumption that $P$ and $p$ are not incident, we obtain the conclusion that the three conics $\mathbf{{R}}_{i}$ are mutually in double contact with chords of contact $\Lambda^{0n}_{ij}$, however these are concurrent in the point $O$, so by Lemma 6 must be in the same double contact pencil which would imply that the 3 axes $\Lambda^{0n}_{ij}$ coincide which we’ve already seen cannot be the case. Therefore $P$ is incident to $p$. ∎
@@
To complete the cube diagram and prove the theorem, we apply Lemma 13 to obtain a unique quadric containing all three conics that is in ring contact with $\mathbf{{S}}_{i}$ along $\mathbf{{R}}^{0}_{i}$ with ring point $P^{0}_{i}$ for $i=1,2,3$.
Using this fact, we now provide a synthetic proof of the 8-quadric theorem, followed by a self-contained coordinate-based proof which includes a proof of Lemma (13) as part of the construction.
Recall that to complete the proof we must show that for each $j\in\{1,2,3\}$, (i) $\mathbf{{T}}^{0}$ is in ring contact to $\mathbf{{S}}_{j}$; and (ii) all 4 ring planes for the face $\mathcal{F}^{0j}_{km}$ have a common axis.
@@
We now show that condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 3, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 3. ∎
Readers who are content with the above proof may skip ahead to Section 5.
@@ 4.4.3 Refinement of Penrose 8-quadric theorem
The above discussion contains a proof of the following refinement of Thm. 8.
Theorem 14.
Given a regular Penrose 7-quadric configuration. Let $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ be the quadric that, by Thm. 8, completes the configuration. Consider the 3 face axes and 3 face spears of the cube faces meeting at the eighth vertex. Then:
@@ 4.4.4 Proof of Extrusion Lemma
We now turn to the promised proof of Lemma 10: Given a regular Penrose 7-conic configuration, it is possible to extrude it to a regular Penrose 7-quadric configuration.
@@ 4.4.5 Proof of the 8-quadric theorem with a common contact element
5 Examples of 8-quadric configurations
The aim of this section is to illustrate the structure of the 8-quadric theorem by constructing several examples step by step and, by slicing them, to obtain all the famous 2D-porisms mentioned in the Introduction. The geometric proof in the next section provides a direct proof valid also for these examples.
@@ 5 Examples of 8-quadric configurations
Consult Figure 7. We start with a sphere (green) as quadric $\mathbf{{S}}_{0}$. Then we choose 3 planes $p_{0}^{i}$ intersecting the sphere in 3 (blue) rings $\mathbf{{R}}_{0}^{i}$ such that each pair of rings meet in a pair of points (white). They form a conical octahedron, featured above in Lemma 13. The corresponding cones touching the sphere along those rings are taken as quadrics $\mathbf{{T}}^{i}$ (yellow). Now for each pair of cones the intersection $\mathbf{{T}}^{i}\cap\mathbf{{T}}^{j}$ consists of two conics (red). These six conics meet three at a time in the 8 yellow points, the common intersection $\mathbf{{T}}^{1}\cap\mathbf{{T}}^{2}\cap\mathbf{{T}}^{3}$ of all the three cones. We call this configuration a conical cube, whereby the yellow points represent the vertices and each red conic represents a pair of opposite edges of the cube.
@@
In the latter case the final quadric $\mathbf{{T}}^{0}$ is regular (directly given by Lemma 13). Since each quadric $\mathbf{{S}}_{i}$ degenerates to a conic, all 3 rings $\mathbf{{R}}_{i}^{k}$ on it also collapse to this conic and each axis $\Lambda_{ij}^{0k}$ is the intersection of two ring planes. In the former case, on the other hand, the final quadric $\mathbf{{T}}^{0}$ is a point pair. This case, however, cannot as easily be subsumed into the logic of section 4.4 since the final 3 axes are identical: $\Lambda^{01}_{23}=\Lambda^{02}_{13}=\Lambda^{03}_{12}$. To picture this in terms of a limit, we consider a sequence of shrinking ellipsoids, the $\mathbf{{S}}_{i}$ getting flatter and $\mathbf{{T}}^{0}$ getting thinner and finally collapsing to a line segment connecting the two points, all planes through either point ultimately being tangent planes enveloping the ellipsoid in the limit.
@@
Figure 8: Left: Three concurrent conics of the conical cube induce a hexagon circumscribing a circle in the plane $p$ through the ring points of the cones and thus proves the Brianchon theorem. Right: In the case of three non-concurrent conics the 6 points split into two triangles. This figure generates a Poncelet configuration.
@@
In the other case the cones do not admit common tangent planes. See Figure 9, right. Now the rings $\mathbf{{R}}_{i}^{0}$ are obtained by intersecting each cone $\mathbf{{S}}_{i}$ with the corresponding plane $p_{i}^{0}$ which is the join of a pair of vertical axes. The final quadric is again guaranteed by Lemma 13. A planar slice through the centers of the spheres shows that 3 pairs of internal tangents of 3 circles are circumscribed to a conic. This theorem seems to be known only for the case, where the final conic degenerates into a point pair, see [Bak36],[Sal17].
@@
Consider 3 planes $p_{0}^{i}$ tangent to a quadric $\mathbf{{S}}_{0}$ at points $P_{0}^{i}$. In order to see the rings $\mathbf{{R}}_{0}^{i}$ as real objects we choose here a regulus as quadric $\mathbf{{S}}_{0}$, thus each ring is a line pair contained in the regulus. Consult Figure 10. This is the case referred to as X-contact in section 3.2, whereby the ring point is incident to the ring plane. We want to picture these line pairs as ‘quadrics’ $\mathbf{{T}}^{i}$, although the situation is more complex. They are in fact double planes when considered point-wise and double points when considered plane-wise, but in an X-contact pencil we can identify them with the line pairs themselves. We can imagine these ‘quadrics’ as the limit of a sequence of reguli all containing the line pair $\mathbf{{S}}_{0}\cap p_{0}^{i}$, as they get flatter and flatter.
@@
The above examples show the following common structure: The three quadrics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ give rise to 6 degenerate quadrics either by pairwise intersections (Examples 5.1 and 5.3) or by pairwise envelopes (Example 5.2). Depending on the choice of such a triplet $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$, one from each pair, the final quadric $\mathbf{{T}}^{0}$ is either regular or a pair of points (or planes). This is summarized in Figure 12.
@@ 6 Algebraic proof of the Penrose theorems
We now turn to a purely algebraic proof of Penrose’s 8-conic theorem. We begin with describing a function $\mathcal{P}(\mathbf{{S}}_{0},\mathbf{p},\mathbf{q},\mathbf{r},a,b,c)$ whose value is a complete Penrose 8-configuration $\{\mathbf{{S}}_{i},\mathbf{{T}}^{j}\}$ whereby:
  • $\mathbf{{S}}_{0}$ is an arbitrary complete conic,
  • $\{\mathbf{p},\mathbf{q},\mathbf{r}\}$ are three arbitrary weighted lines in $\mathbb{R}P^{2}$, and
  • $\{a,b,c\}$ are real parameters.
Note that the parameter space of $\mathcal{P}$ is 17-dimensional: $\mathbf{{S}}_{0}$ depends on 5 parameters, each of the weighted lines $\{\mathbf{p},\mathbf{q},\mathbf{r}\}$ depends on 3 real parameters, and $\{a,b,c\}$ contributes 3 more real parameters for a total of 17.
The following discussion presents and verifies the explicit formulas provided by the algorithm for the cube $\{\mathbf{{S}}_{i},\mathbf{{T}}^{j}\}$. It makes repeated use of the fact that two point-wise conics are in double contact with each other if and only if their pencil contains a double line. That is, there exists $a$ and $b$ such that
@@
where ${m}$ is the joining line of the contact points. This equation can obtained from Eq. 6 by translating the latter into abbreviated notation (Section 3.3.1). Note that $\mathbf{m}^{2}$ has to be non-zero to establish double contact, so that we can normalize $a$ and $b$ to have $\mathbf{m}^{2}$ alone on the right-hand side.
It focuses on the case of point-wise conics; similar formulas exist also for line-wise conics.
6.1 The Penrose formulas
We start with an arbitrary conic $\mathbf{{S}}_{0}$ and a line $\mathbf{p}$. Consider the sum:
@@ 6.1 The Penrose formulas
Then $\mathbf{{T}}^{1}$ and $\mathbf{{S}}_{0}$ are in double contact because we have $\mathbf{{T}}^{1}-\mathbf{{S}}_{0}=\mathbf{p}^{2}$ as in Eq. 45.
@@
As discussed in Section 4.2, the conic $\mathbf{{S}}_{0}$ can move along two conics in double contact with it, while staying in double contact with both. The family of such conics can be expressed using $\mathbf{{S}}_{0}$, the chords of contact $\mathbf{q}$ and $\mathbf{r}$, and a free parameter $a$:
@@
6.2 The algebraic proof
The algebraic proof of the 8-conic theorem reverses this process. Given a Penrose 7-configuration, it calculates arguments $(\mathbf{p},\mathbf{q},\mathbf{r},a,b,c)$ which produces this configuration, and uses them, along with $\mathbf{{S}}_{0}$, to calculate $\mathbf{{T}}^{0}$.
Let $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{j}$ be given as a Penrose 7-configuration, without $\mathbf{{T}}^{0}$. With suitable normalization of $\mathbf{{T}}_{i}$, we can apply Eq. 46 to obtain:
$\mathbf{p}^{2}=\mathbf{{S}}_{0}-\mathbf{{T}}^{1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{q}^{2}=\mathbf{{S}}_{0}-\mathbf{{T}}^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{r}^{2}=\mathbf{{S}}_{0}-\mathbf{{T}}^{3}$ (51)
Each of the real parameters $\{a,b,c\}$ corresponds to a complete face of the given Penrose 7-configuration. For example, $a$ corresponds to face $\mathcal{F}^{23}_{01}$, whose chords of contact pass through $P^{23}_{01}$. We seek $a$ so that $\mathbf{{S}}_{1}$ satisfies Eq. 47. By Section 4.2, this value can be directly obtained in terms of $\mathbf{{S}}_{0},\mathbf{{q}},$ and $\mathbf{{r}}$. Similarly for $b$ and $c$. Finally, we apply Eq. 6.1 to obtain the desired eighth conic $\mathbf{{T}}^{0}$, completing the cube and the proof.